# Push and Pull: A Framework for Measuring Attentional Agency

### v1

@@ 3.1. General Setup and Key Definitions
We study a canonical interaction between three parties: an agent, who wants to pay attention to informational objects they themselves value highly; an advocate, who also has preferences about what the agent attends to; and a platform that determines whether and how informational objects enter the agent’s attentional field.
The agent has type $\theta$ drawn from a distribution $F\in\Delta(\Theta)$ where $\Theta$ is a finite type space. There is a universe of informational objects represented by an ordered set $X=(x_{1},x_{2},\dots)$. An allocation is a bijective function $\phi:\mathbb{N}\to X$. The agent has a type-dependent value for each informational object that is represented by a utility function $u:X\times\Theta\to\mathbb{R}_{\geq 0}$. Owing to their limited attention, the agent discounts informational objects based on how they are allocated according to a weakly decreasing function $\delta:\mathbb{N}\to[0,1]$ where $\delta_{0}=1$ without loss of generality. The agent’s value of allocation $\phi$ is
@@
The platform supplies the agent with an allocation. Let $\Gamma$ denote a partition of $X$, which induces what we will call a allocation technology. [^1] A partition of $X$ is a set of subsets of $X$ (“partitions” or “blocks”) such that every element of $X$ belong to one and only one block. Specifically, we assume that the platform can produce allocations that rearrange the order of the partition blocks of $\Gamma$, but not move objects within blocks. The interpretation is that digital information is naturally bundled together into objects (e.g., web pages, academic articles, social media posts), and greater degrees of technological sophistication are required to break apart, comprehend, extrapolate preferences to, and recombine more granular units of information within these objects. Let $\Phi(\Gamma)=\{\phi:\mathbb{N}\to X\,|\,\forall\gamma\in\Gamma,\text{ if }x_{i},x_{j}\in\gamma\text{ then }\phi^{-1}(x_{i})-\phi^{-1}(x_{j})=i-j\}$ denote the set of all possible allocations that can be constructed by permuting the blocks of $\Gamma$.
@@
for a given weight $\lambda\in[0,1]$. For simplicity, we will assume throughout that the primitives of the model are such that the solution to this problem exists and is unique (or can be arbitrarily chosen from a set of candidate solutions; this assumption does not materially affect the points we raise in our discussion). Next, define
@@
Pull is defined as the amount of informational value a platform delivers the agent as a fraction of the total value it is technologically capable of delivering them. Formally, for a given $u,v,\lambda$, $\Gamma$, and $m$, a platform’s pull is
@@
Push and Pull both range from 00 to $1$ and index the amount of attentional value that is being delivered to agents and advocates, respectively. According to these definition, it is always the case that increasing push requires that one (weakly) decrease pull. Note, however, that push and pull are not necessarily zero-sum. We analyze the conditions that determine the relative tradeoffs between push and pull fully in the next section.
Another important point is that our framework defines push and pull for a single agent, advocate, and platform interaction. Across many interactions, each platform produces a distribution of Pull and Push values which could be analyzed in terms of their mean, variance, group-level disparities, and other distributional properties.
3.2. Analysis and Discussion
In this section we highlight some of its most salient aspects of the above framework and definitions.
@@ 3.2. Analysis and Discussion
First, note that $U_{\lambda}$ is a weakly decreasing function of $\lambda$ and $V_{\lambda}$ is a weakly increasing function of $\lambda$ [^2] This is a direct consequence of the envelope theorem. However, the rate at which these functions change depends upon the relative alignment of agent and advocate preferences. Taking the platform’s allocation technology and agent model as given, we can consider the frontier of feasible push and pull values as a function of these underlying preferences. The platform effectively selects a point on this frontier by setting $\lambda$.
In a “zero-sum” situation where $u=-v$, then the platform’s objective simplifies to $(2\lambda-1)\,\mathbb{E}[U(\phi,\theta)|m]$, and they must choose either to deliver the agent’s most-preferred allocation (if $\lambda>\frac{1}{2}$) or their least-preferred allocation (if $\lambda<\frac{1}{2}$). In such an interaction, it is not only the case that pull and push are themselves zero sum ($\textbf{Pull}+\textbf{Push}=1$), but also that they change sharply around a critical value of $\lambda$. In other words, when agents and advocate push and pull in opposite directions, the platform will only accord agency to one of them. Conversely, if preferences are fully coincidental ($u=v$), then the platform can achieve maximal pull and push simultaneously ($\textbf{Pull}=\textbf{Push}=1$) for any value of $\lambda$ because the same allocation is optimal for both agents.
A less intuitive case of simultaneously maximal pull and push occurs when the agent is indifferent over certain aspects of their allocation that the advocate has strict preferences over and vice-versa. When preferences are “orthogonal” in this way, the platform can also achieve maximal pull and push simultaneously ($\textbf{Pull}=\textbf{Push}=1$) for any interior value of $\lambda$. From a policy perspective, situations of this variety have been described are ideal targets for “nudge”-style interventions (Thaler and Sunstein, 2009), given that they can, in theory, achieve a planner’s objectives without decreasing people’s attentional agency.
@@ 3.2.2. Model advances increase platform value, but also redistribute attentional agency
Advances in machine learning and artificial intelligence (e.g., new architectures, more parameters, or additional compute (Kaplan et al., 2020)) enter our framework in two ways. The first is through $m$, which represents the depth and specificity with which the platform is able to infer an agent’s type $\theta$ before constructing an informational allocation.
Critically, however, such changes enhance the platform’s ability to provide value to both agent and advocate. In the former case, accurate inferences about an agent’s type enables the platform to more successfully extrapolate their preferences to hypothetical allocations. Simultaneously, these inferences also enable the platform to facilitate more precise targeting by advocates. Indeed, recent work shows that frontier models are capable of greater persuasion precisely because they are better at “matching the language or content of a message to the psychological profile of its recipient” (Matz et al., 2024), increasing the stakes of strategic misalignment between agents and advocates.
The net effect of model improvements on attentional agency therefore depend on whether the innovations are more useful for predicting marginal differences in $u$ or $v$ across various allocations. Recall that the platform’s objective when selecting a allocation is to maximize the linear combination
@@
Although the net effect on pull and push will depend, in each case, on which of these effects prevails, one thing is clear: the stakes of informational agency only grow as models become more sophisticated.
@@ 3.2.3. AI enables attentional allocation to operate over more granular and varied information objects
Multiple features of modern generative foundational models (Bommasani et al., 2021) now enable AI systems to achieve unprecedented levels of informational granularity: (1) attentional transformers parse a wide variety of data sources into their constitutive tokens, meaning that systems can apprehended complex informational objects at the scale of individual symbols (Vaswani et al., 2017); (2) ever-expanding context windows enable sequence-based models to capture long-range data dependencies (Press et al., 2021; Su et al., 2024); (3) multi-modality is in the process of extending the informational purview of AI systems to text, image, video, audio, and other formats (Yin et al., 2023); and (4) generative capabilities enable systems to flexibly recombine these informational atoms into arbitrary new structures that do not necessarily resemble the sources they were initially drawn from. Although it has always been possible for platforms to disassemble informational objects in a literal sense, foundation models enable a categorically greater degree of sophistication when extrapolating an agent’s preferences and intentions to these smaller units of information.
@@
Within our framework, the cumulative effect of these innovations is to refine the platform’s informational domain into a finer partition $\Gamma^{\prime}\subset\Gamma$.  [^4] Specifically, if for every $\gamma^{\prime}\in\Gamma^{\prime}$ there exists a $\gamma\in\Gamma$ such that $\gamma^{\prime}\subseteq\gamma$ in the sense of weak subsetting, with at least one subset relationship strict. The additional degrees of freedom this opens up expand the space of possible allocations and, consequently, enable the platform to deliver more value overall. Formally, $\Phi(\Gamma)\subset\Phi(\Gamma^{\prime})$ so that ${P^{\prime}}_{\lambda}\geq P_{\lambda}$ for any given $\lambda$. As in the case of model development, the ultimate impact of such a change on welfare and informational agency will depend on the relative alignment of preferences and other primitives of the model.
@@
One clear take-away, however, is that highly capable multi-modal AI models which are fully aligned to agents (i.e., for which $\lambda=1$ so that $\textbf{Pull}=1$) increase their ability to extract informational valuable from a variety of sources without satisfying the implicit “attentional bargain” that underlies many aspects of the internet’s current economic model—branding, sponsored content, advertising, and other features that businesses employ to monetize their content. As (Wu, 2017) points out in his history of the attention economy, the basic strategy to “draw attention with apparently free stuff and then resell it” has long been a part of the advertising business. Technological innovations intermittently disrupt the status quo implementation of this strategy, as exemplified by the famous “Betamax case” over whether Videocassette recordings of television shows (which enabled people to re-watch them without embedded advertisements) constituted copyright infringement. [^5] Sony Corp. of America v. Universal City Studios, Inc., 1984.
@@
In contrast to prior waves of creative destruction, however, the very generality of general artificial intelligence means that one technology stands to disrupt nearly every existing informational format simultaneously—a development precedented only, perhaps, by advent of digital computers. The advent of human or super-human level artificial intelligence may represent a final frontier in these debates: by effectively solving the attention problem, these technology forces us to confront the incentive problem head on. We return to the question of how these insights inform policy approaches to refactoring the landscape of attentional agency in Section 5.3, below.
4. Attentional Agency in Present and Future Platforms
In the first part of this section, we describe how our framework maps on to five existing categories of internet platform: search, content libraries, social media, matching, and marketplaces. In the second part of this section, we apply our framework to consider the impact that advances in artificial intelligence are having on these platforms.
@@ 4. Attentional Agency in Present and Future Platforms
Table 1. How prominent categories of informational platforms map on to our formal framework.
@@
Table 1 provides a high level overview of how various categories of platform map on to our framework. A side-by-side comparison reveals not only that many platforms are fundamentally oriented around attentional allocation, but that the boundaries separating these categories are becoming increasingly blurred over time. Search engines that may have once depended purely on a user’s entered query and a few contextual variables (e.g. location) now incorporate user history (Matthijs and Radlinski, 2011). Social media platforms often emphasize a feed (with minimal interaction options beyond “show me more content”) but may also support search. Online marketplaces sell a mixture of physical and digital goods, but now often offer recommendation features.
Below, we briefly highlight key characteristics of these five categories, then note how our model helps to clarify their conceptual similarities and differences. We then return to the topic of generative foundation models, which are beginning to blur some of the distinctions that have traditionally separated platform categories (Deldjoo et al., 2024).
4.1. Mapping Platform Categories Onto the Framework
@@
Search Engine: Search engines rank web pages based on user queries (Varian, 2016; Resnick and Varian, 1997). Agents engaged in search generally want to specify their query as briefly as possible, meaning that search engines are most efficient when they can anticipate people’s needs and desires. This tends to incentivize platforms to build up a deep understanding of preferences and intention through repeated interaction with individual users. Interactions with the platform (e.g., click-through rates) are also used to tailor future results delivered to the same or other users (Resnick and Varian, 1997).
A central challenge of general web search is that pages are heterogeneous and constantly evolving, meaning that one must continuously crawl, index, and calculate the relevance of a very large number of informational objects to maintain relevance. As PageRank proved early on, the network structure of the web itself can be used as a proxy for relevance. Increasing levels of algorithmic sophistication have enabled search providers to predict the general quality and specific relevance of web pages with greater degrees of sophistication over time.
@@
Content Library: Content libraries provide access to large collections of media and generally enable users to both search for items they already know and browse recommendations to discover new items. Media preferences tend to correlate significantly across individuals, meaning that general popularity is often a strong proxy for relevance to individual users. Personalization can significantly enhance the value of recommendations by inferring people’s particular tastes (e.g., a love of horror films) (Gomez-Uribe and Hunt, 2015).
@@
Social Media: Recommendation-driven content feeds, such as those found on social media platforms, present users with a curated stream of multimedia objects centered around the activity of other users. Many social media feeds orient the user’s attention around a single activity: requesting more content. Social media platforms also tend to primarily infer an agent’s interests and intentions implicitly (e.g., from dwell time), rather than explicitly (e.g, entering a query). Social media is also distinguished by the fact that most individuals who act as agents also act as advocates: people both want to “see and be seen.”
Matching: Preferences on matching platforms are highly specific and driven by a clear objective, such as finding a compatible partner or securing a desired job. Matching platforms are also characterized by the fact that many users act as both agents and advocates. For example, successfully matching on a dating app requires other users to view and like one’s profile. Many matching platforms allow advocates (premium users) to gain visibility by influencing the attention of other users.
@@
Summary: Although matching platforms are relatively easy to identify (items being ranked are platform participants themselves), marketplaces can be characterized by the actual exchange of consumer goods, and general search engines are relatively clearly defined, both content libraries and social media often offer a mixture of search, top-n recommendation, and infinite scroll content feeds.
Nonetheless, when we view each of these platforms as performing the core operation of delivery users ranked lists of some kind of item (be that web pages, videos, posts, other users, or offerings of physical goods), any of them can lean more towards pushing or pulling, and this decision is likely to be made on the basis of what is most profitable for that platform (barring policy interventions, such as those we discuss below).
4.2. The Effect of Generative Foundation Models
Table 2 summarizes how AI search tools (powered by generative foundation models) compare to the five existing platform categories just reviewed. As discussed in Section 3, these systems operate at the level of individual tokens (Bommasani et al., 2021), which enables them to flexibly tailor content to the user’s specific preferences ($\theta$), as inferred from their interaction history and context. The capacity of modern AI systems to draw meaningful connections between distal information sources and modalities (see Section 3.2.3) strengthens complementarities between various data sources: scientific papers, law articles, encyclopedia entries, code, and other sources of knowledge that were previously mostly referenced independently can now be interrelated in new and more valuable ways.

### v2

@@ 3.1. General Setup and Key Definitions
We study a canonical interaction between three parties: an agent, who wants to pay attention to informational objects they themselves value highly; an advocate, who also has preferences about what the agent attends to; and a platform that determines whether and how informational objects enter the agent’s attention.
The agent has type $\theta$ drawn from a distribution $F\in\Delta(\Theta)$ where $\Theta$ is a finite type space. There is a universe of informational objects represented by an ordered set $X=(x_{1},x_{2},\dots)$. An allocation is a bijective function $\phi:\mathbb{N}\to X$.
The agent has a type-dependent value for each informational object that is represented by a utility function $u:X\times\Theta\to\mathbb{R}_{\geq 0}$. Owing to their limited attention, the agent discounts informational objects based on how they are allocated according to a weakly decreasing function $\delta:\mathbb{N}\to[0,1]$, where $\delta_{0}=1$ without loss of generality. The agent’s value of allocation $\phi$ is
@@
The platform supplies the agent with an allocation. Let $\Gamma$ denote a partition of $X$, which induces what we will call an allocation technology. [^3] A partition of $X$ is a set of subsets of $X$ (“partitions” or “blocks”) such that every element of $X$ belongs to one and only one block. Specifically, we assume that the platform can produce allocations that rearrange the order of the partition blocks of $\Gamma$, but not move objects within blocks. The interpretation is that digital information is naturally bundled together into objects (e.g., web pages, academic articles, social media posts), and greater degrees of technological sophistication are required to break apart, comprehend, extrapolate preferences to, and recombine more granular units of information within these objects. Let $\Phi(\Gamma)=\{\phi:\mathbb{N}\to X\,|\,\forall\gamma\in\Gamma,\text{ if }x_{i},x_{j}\in\gamma\text{ then }\phi^{-1}(x_{i})-\phi^{-1}(x_{j})=i-j\}$ denote the set of all possible allocations that can be constructed by permuting the blocks of $\Gamma$.
@@
for a given weight $\lambda\in[0,1]$. [^4] In general, the parameter $\lambda$ is a choice variable under the control of the platforms. However, even if platforms can freely select $\lambda$ in principle, the full range of choices may not be feasible in practice. Platforms operate within a broader “attention economy” and consequently must compete with other activities for user engagement (most notably, the services provided by competitors). Market forces therefore constrain the minimum and maximum amount of $\lambda$ a platform can select without losing agent engagement or advocate buy-in entirely. [^5] An important exception occurs when platforms intentionally withhold attentional value to induce users to behave in a certain way—for example, when a dating app restricts how many profiles a user can view without paying for a premium version of the service. We note, however, that even in such cases, platforms typically incentivize behaviors by promising to move users onto or along the optimal push-pull frontier defined by the maximization (3). For simplicity, we will assume throughout that the primitives of the model are such that the solution to this problem exists and is unique (or can be arbitrarily chosen from a set of candidate solutions; this assumption does not materially affect the points we raise in our discussion). Next, define
@@
Pull is the amount of informational value a platform delivers the agent as a fraction of the total value it is technologically capable of delivering them. Formally, for a given $u,v,\lambda$, $\Gamma$, and $m$, a platform’s pull is
@@
Push and Pull both range from 00 to $1$ and index the amount of attentional value that is being delivered to agents and advocates, respectively. According to these definitions, it is always the case that increasing push (weakly) decreases pull. Note, however, that push and pull are not necessarily zero-sum. We analyze the conditions that determine the relative tradeoffs between push and pull fully in the next section.
Another important point is that our framework defines push and pull for a single agent, advocate, and platform interaction. Across many interactions, each platform produces a distribution of Pull and Push values which could be analyzed in terms of their mean, variance, group-level differences, and other distributional properties.
3.2. Analysis
In this section we highlight a few of the framework’s most salient features.
@@ 3.2. Analysis
First, note that $U_{\lambda}$ is a weakly decreasing function of $\lambda$ and $V_{\lambda}$ is a weakly increasing function of $\lambda$ [^6] This is a direct consequence of the envelope theorem. See (Carter, 2001), pg. 603. for a formal statement and introduction to the theorem. However, the rate at which these functions change depends upon the relative alignment of agent and advocate preferences. Taking the platform’s allocation technology and agent model as given, we can consider the frontier of feasible push and pull values as a function of these underlying preferences. The platform effectively selects a point on this frontier by setting $\lambda$.
In a “zero-sum” situation where $u=-v$, the platform’s objective simplifies to $(2\lambda-1)\,\mathbb{E}[U(\phi,\theta)|m]$, and they must choose either to deliver the agent’s most-preferred allocation (if $\lambda>\frac{1}{2}$) or their least-preferred allocation (if $\lambda<\frac{1}{2}$). In such an interaction, it is not only the case that pull and push are zero sum ($\textbf{Pull}+\textbf{Push}=1$), but also that they change sharply around a critical value of $\lambda$. In other words, when agents and advocates push and pull in opposite directions, the platform will only accord agency to one of them. Conversely, if preferences are fully coincidental ($u=v$), then the platform can achieve maximal pull and push simultaneously ($\textbf{Pull}=\textbf{Push}=1$) for any value of $\lambda$ because the same allocation is optimal for both agents.
A less intuitive case of simultaneously maximal pull and push occurs when the agent is indifferent over certain aspects of their allocation that the advocate has strict preferences over and vice-versa. When preferences are “orthogonal” in this way, the platform can also achieve maximal pull and push simultaneously ($\textbf{Pull}=\textbf{Push}=1$) for any interior value of $\lambda$. From a policy perspective, situations of this variety are ideal targets for “nudges” (Thaler and Sunstein, 2009), given that they can, in theory, achieve a planner’s objectives without decreasing people’s attentional agency.
@@ 3.2.2. Model advances increase platform value, but also redistribute attentional agency
Advances in machine learning and artificial intelligence (e.g., new architectures, more parameters, or additional compute (Kaplan et al., 2020)) enter our framework in two ways. The first is through $m$, which represents the depth and specificity with which the platform infers an agent’s type $\theta$ before constructing an informational allocation.  [^8] The generality of this framework means that it can be applied to any recommender system, regardless of the specific algorithm used. See, e.g., (Cai and Daskalakis, 2022) for detailed treatment of how a seller might learn agents’ types in a mechanism design context and (McAuley, 2022) for a general overview of user modeling.
Critically, however, such changes enhance the platform’s ability to provide value to both agent and advocate. In the former case, accurate inferences about an agent’s type enable the platform to more successfully extrapolate their preferences to hypothetical allocations. Simultaneously, these inferences also enable the platform to facilitate more precise targeting by advocates. Indeed, recent work shows that frontier models are capable of greater persuasion precisely because they are better at “matching the language or content of a message to the psychological profile of its recipient” (Matz et al., 2024), increasing the stakes of strategic misalignment between agents and advocates.
The net effect of model improvements on attentional agency therefore depends on whether the innovations are more useful for predicting marginal differences in $u$ or $v$ across various allocations. Recall that the platform’s objective when selecting an allocation is to maximize the linear combination
@@
Although the net effect on pull and push will depend, in each case, on which of these effects prevails, one implication always holds: the stakes of informational agency consistently grow as models become more sophisticated.
@@ 3.2.3. AI enables attentional allocation to operate over more granular and varied information objects
Multiple features of modern generative foundational models (Bommasani et al., 2021) now enable AI systems to achieve unprecedented levels of informational granularity: (1) attentional transformers parse a wide variety of data sources into their constitutive tokens, meaning that systems can apprehend complex informational objects at the scale of individual symbols (Vaswani et al., 2017); (2) expanded context windows enable sequence-based models to capture long-range data dependencies (Press et al., 2021; Su et al., 2024); (3) multi-modality is in the process of extending the informational purview of AI systems to text, image, video, audio, and other formats (Yin et al., 2023); and (4) generative capabilities enable systems to flexibly recombine these informational atoms into arbitrary new structures that do not necessarily resemble the sources they were initially drawn from. Although it has always been possible for platforms to disassemble informational objects in a literal sense, foundation models enable a categorically greater degree of sophistication when extrapolating an agent’s preferences and intentions to these smaller units of information.  [^9] There may be cases in which an agent’s utility function $u$ contains a preference for less granular partitions. For instance, consider an agent that prefers to listen to music only in the form of complete and correctly ordered albums.
@@
Within our framework, the cumulative effect of these innovations is to refine the platform’s informational domain into a finer partition $\Gamma^{\prime}\subset\Gamma$.  [^10] Specifically, if for every $\gamma^{\prime}\in\Gamma^{\prime}$ there exists a $\gamma\in\Gamma$ such that $\gamma^{\prime}\subseteq\gamma$ in the sense of weak subsetting, where at least one inclusion is strict. The additional degrees of freedom this opens up expand the space of possible allocations and, consequently, enable the platform to deliver more value overall. Formally, $\Phi(\Gamma)\subset\Phi(\Gamma^{\prime})$ so that ${P^{\prime}}_{\lambda}\geq P_{\lambda}$ for any given $\lambda$. As in the case of model development, the ultimate impact of such a change on welfare and informational agency will depend on the relative alignment of preferences and other primitives of the model.
@@
One clear take-away, however, is that highly capable multi-modal AI models which are fully aligned to agents (i.e., for which $\lambda=1$ so that $\textbf{Pull}=1$) increase their ability to extract informational value from a variety of sources without satisfying the implicit “attentional bargain” that supports much of the internet’s current economic model—branding, sponsored content, advertising, and other features that businesses employ to monetize their content. As (Wu, 2017) points out in his history of the attention economy, the basic strategy to “draw attention with apparently free stuff and then resell it” has long been a part of the advertising business. Technological innovations intermittently disrupt the status quo implementation of this strategy, as exemplified by the famous “Betamax case” over whether Videocassette recordings of television shows (which enabled people to re-watch them without embedded advertisements) constituted copyright infringement. [^11] Sony Corp. of America v. Universal City Studios, Inc., 1984.
@@
In contrast to other drivers of creative destruction, however, the very generality of general artificial intelligence means that a single technology stands to disrupt nearly every existing informational format simultaneously. The advent of human or super-human level artificial intelligence may represent a final frontier in these debates, effectively solving the attention problem. We return to the question of how these insights inform policy approaches to refactoring the landscape of attentional agency in Section 5.4, below.
4. Attentional Agency on Platforms
@@
Table 1. How prominent digital platform categories map onto our framework.
[…]
How prominent digital platform categories map onto our framework.
@@
Table 1 provides a high-level overview of how platform categories map on to our framework. Comparing various platforms side-by-side highlights not only their common approach to attentional allocation, but also that the boundaries separating these categories are becoming increasingly blurred over time. Search engines that may have once depended purely on a user’s entered query and a few contextual variables (e.g. location) now incorporate user history (Matthijs and Radlinski, 2011). Social media platforms often emphasize a feed (with minimal interaction options beyond “show me more content”) but may also support search. Online marketplaces sell a mixture of physical and digital goods, but now often offer recommendation features.
In what follows, we briefly review how our framework maps onto existing platform categories, then briefly discuss how generative foundation models are further eroding the distinctions that have traditionally separated them (Deldjoo et al., 2024).
@@
Web Search Engine: Search engines primarily rank web pages based on user queries and intrinsic, site-specific features (Varian, 2016; Resnick and Varian, 1997). Agents engaged in web search generally want to specify their query as briefly as possible, meaning that search engines are most efficient when they can anticipate people’s needs and desires (Resnick and Varian, 1997; Matthijs and Radlinski, 2011). This tends to incentivize platforms to build up a deep understanding of preferences and intention through repeated interaction with each individual user. One user’s interactions with a platform are also used to tailor results delivered to similar users through techniques such as collaborative filtering.
Queries enable users to locate relevant content efficiently, but they also equip platforms with highly specific data about user characteristics and momentary intentions. Consequently, many leading search engines harness this information by allowing web pages to augment their ranking preferentially among users with specific characteristics.
@@
Content Library: Content libraries provide access to large collections of media and generally enable users to both search for items they already know and browse recommendations to discover new items. Media preferences tend to correlate significantly across and within individuals, meaning that both general popularity and personalization can significantly enhance the value of recommendations (Gomez-Uribe and Hunt, 2015).
In the context of content libraries, pull could be decreased because of platform promotion (e.g., a video streaming service promoting a show they funded), but also through limitations of collaborative filtering.
@@
Social Media: Social media platforms present users with a curated stream of multimedia objects centered around the activities of other users. Many social media feeds orient the user’s attention around a single activity: requesting more content. Social media platforms tend to primarily infer an agent’s interests and intentions implicitly (e.g., from dwell time), rather than explicitly (e.g, entering a query). Social media is also distinguished by the fact that many users want to both “see and be seen,” acting as both agents and advocates.
Matching: Matching platforms help interested parties find one another. Some such platforms, such as dating apps, are “two-sided,” with users acting as both agents and advocates. Many matching platforms allow advocates (premium users) to gain visibility by influencing the attention of other users.
@@
AI Search: Table 2 summarizes how AI search tools (powered by generative foundation models) compare to the five existing platform categories just reviewed. As discussed in Section 3, AI search tools operate at the level of individual tokens (Bommasani et al., 2021), which enables them to flexibly tailor content to the user’s specific preferences ($\theta$), as inferred from their interaction history and context. The capacity of modern AI systems to draw meaningful connections between distal information sources and modalities (see Section 3.2.3) strengthens complementarities between various data sources: scientific papers, law articles, encyclopedia entries, code, and other sources of knowledge that were previously mostly referenced independently can now be interrelated in new and more valuable ways.
@@
[…]
Applying the push/pull framework to AI search, chat, and assistant tools, as powered by modern generative foundation models.